OpenStudy (anonymous):

This year, a 30-second commercial to be shown during the Super Bowl will cost $4 million. How much would a 20-second commercial cost? A 40-second commercial? An estimated 110 million people will be watching this year. About how much money is being spent per person in a 30-second ad?

OpenStudy (anonymous):

I cannot figure out how to set it up is it a ratio and inequality or equation just started these in 6th grade.

OpenStudy (lucaz):

maybe you can use 4mi/30s = xmi/20s

OpenStudy (anonymous):

does that mean 4 million divided by 30 seconds =x millionlsdivided by 20 seconds

OpenStudy (lucaz):

yes, just a proportion

OpenStudy (anonymous):

I know how to do the secohd part then multiply the answer for 20sec by 2 to get the cost for 40 second commercail is that right

OpenStudy (lucaz):

the second part you mean "About how much money is being spent per person in a 30-second ad?"

OpenStudy (anonymous):

sorry I meant how much is a 40 second ad

OpenStudy (lucaz):

oh, well you can use the same process, 4mi/30s = xmi/40s now for 40s

OpenStudy (anonymous):

The 4 million would be diveded by the people viewng for the amount spent for the 30 second add

OpenStudy (anonymous):

oh thank you

OpenStudy (lucaz):

this las question I don't get it =/ not sure how to answer it, sorry

OpenStudy (anonymous):

The last part of the question is is C people watch the commercial then how much will be spen on each person. It gives the amount of people who will watch and we know the commercial cost 4 m illion so I will divide the people by the 4 million

OpenStudy (anonymous):

This year, a 30-second commercial to be shown during the Super Bowl will cost $4 million. An estimated 110 million people will be watching this year. About how much money is being spent per person in a 30-second ad?

OpenStudy (lucaz):

I think it's something like this 1 person whatching costs,let's say, $3 2 people = $6 3 people = $9 you can divide the cost of a 30s commercial by 110 mi people

OpenStudy (lucaz):

this gives you the money per person..

OpenStudy (anonymous):

4,000,000/110.000.000=x

OpenStudy (lucaz):

if we're doing it correctly thats how you find the answer

OpenStudy (anonymous):

Thanks I just needed someone to help me set up and to agree I really apprecuated your help.

OpenStudy (lucaz):

you're welcome..